# UniAttn: Reducing Inference Costs via Softmax Unification for Post-Training LLMs

### v1

@@ 3.2 Analysis on Existing Methods
$\mathbf{x}_{i+1}=\mathbf{x}_{i}+\delta,\quad\text{where}\quad||\delta||\ll||\mathbf{x}_{i}||.$ (6)
We define $\delta$ as the “depth factor” that model applies on activations for further discussion. $\delta$ represents the extent to which model depth influences activations.
Analysis on Cross-layer KV Sharing: We adopt 3.5 to analyze the cross-layer KV sharing architecture, in which KV-cache is shared across multiple consecutive layers. Suppose layer $i+1$ shares the KV-cache from layer $i$, the MHA operation in layer $i+1$ can be written as:
$\mathbf{x}^{\prime}_{i+1}=\text{softmax}(\frac{QK^{T}_{i}}{\sqrt{d_{k}}})V_{i}W_{o}+\mathbf{x}_{i+1},$ (7)
@@
Proof.
$\mathbf{x}^{\prime}_{i+1}$ $=\text{softmax}(\frac{\mathbf{x}_{i+1}W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}})V_{i}W_{o}+\mathbf{x}_{i+1}$ (8)
$=\text{softmax}(\frac{(\mathbf{x}_{i}+\delta)W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}})V_{i}W_{o}+\mathbf{x}_{i+1}$
$=\text{softmax}(\frac{\mathbf{x}_{i}W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}}+\frac{\delta W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}})V_{i}W_{o}+\mathbf{x}_{i+1}$
Let $A_{i}=\frac{\mathbf{x}_{i}W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}}$, $\hat{\delta}=\frac{\delta W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}}$. By expanding Softmax to the first order:
$\mathbf{x}^{\prime}_{i+1}$ $=\big{[}\text{softmax}(A_{i})+J(A_{i})\hat{\delta}\big{]}V_{i}W_{o}+\mathbf{x}_{i+1}+o(\hat{\delta}^{2}),$ (9)
where $J(A_{i})$ is the Jacobian of Softmax at $A_{i}$. We ignore the $o(\hat{\delta}^{2})$ remainder. (Xiao et al., 2024) has discovered that the pre-trained LLMs generally exhibit an “attention sink” feature, in which self-attentions heavily attend to the “sink token” at the beginning of a sequence. We adopt such feature to estimate $||J(A_{i})||$. According to the statistical pattern concluded by (Xiao et al., 2024), attention logits $\mathbf{a}\in\mathbb{R}^{l}$ of a sequence with length $l$ is approximately:
$a_{i}=1\;(0<i<d),\;a_{j}=-1\;(i\geq d),\;d\ll l$ (10)
(Xiao et al., 2024) empirically adopts $d=4$. For a sequence of $l=1024$, $||J(\text{softmax}(\mathbf{a}))||\approx 0.03\ll 1$, and decreases as $l$ increases. Therefore, we can conclude that the coefficient of the $\hat{\delta}$ term is significant smaller than that of the Softmax term, making the contribution of the “depth factor” negligible. ∎
Proposition 3.6 shows that cross-layer KV sharing diminishes the impact of model depth on model activations. When further analyzing the Softmax term, we derive the following:
$\mathbf{x}^{\prime}_{i+1}$ $=\text{softmax}(A_{i})V_{i}W_{o}+\mathbf{x}_{i+1}$ (11)
$=\text{softmax}(\frac{\mathbf{x}_{i}W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}})V_{i}W_{o}+\mathbf{x}_{i+1}.$
The only difference between the first term in Equation 11 and conducting MHA on $\mathbf{x}^{\prime}_{i}$ is that they use different $W_{q}$ matrices. Consequently, the MHA module in layer $i+1$ can be interpreted as another query head for the MHA module in layer $i$. In essence, cross-layer KV sharing transforms MHA modules in different layers into additional query heads in a single layer.
Figure 2: Cosine similarity results of average Softmax activations. Across all settings, the average Softmax activations of the top half of layers share a high cosine similarity very close to 1.
3.3 UniAttn: Softmax Unification in Attention
Figure 3: Pipeline comparison between standard decoder-based transformers, CLA (Brandon et al., 2024), and UniAttn. UniAttn shares the Softmax activations across layers in grouped Superblocks and adds linear transformation $W_{c}$ to compensate for the unification error.
To achieve both memory-efficiency and inference acceleration, we investigate the Softmax operation because (1) calculating Softmax demands the entire K-cache that accounts for 50% memory of the KV-cache; and (2) multiple studies (Geng et al., 2018; Koohpayegani & Pirsiavash, 2024) have recognized that the Softmax operation leads to high inference latency. We study the redundancy of Softmax activations in pre-trained LLMs by measuring the importance of Softmax operations in each layer. To ensure the generalizability of our study, we employ post-training datasets PMC (Wu et al., 2023), Infinity Instruct (Math) (BAAI, 2024), Tulu3 (Lambert et al., 2024), and long samples (F15K) provided by (Su, 2023), and utilize LLaMA-2 7B (Touvron et al., 2023), LLaMA-3.1 8B (Dubey et al., 2024), Mistral 7B (Jiang et al., 2023), and Gemma-2 9B (Rivière et al., 2024). We collect all Softmax activations, i.e., the $s$ vectors in Equation 1, and compute the cosine similarity $\text{Sim}(i)=\cos(s_{i},s_{i-1})$ for each sequence. We then average $\text{Sim}(i)$ over all sequences. As shown in Figure 2, Softmax activations in top half layers share a high cross-layer similarity across different experiment settings. This observation shows that Softmax operations are highly redundant in LLMs.
Motivated by such redundancy, we propose Softmax Unification in Attention (UniAttn) for post-training LLMs. Specifically, UniAttn groups consecutive decoders in LLMs as “Superblocks”, and unifies the Softmax activations inside each Superblock to reduce the full Softmax operations and KV-cache size during inference. Additionally, we observe that the error introduced by Softmax unification can be compensated by a linear projection, and in turn design a pipeline to initialize and train the linear projection. UniAttn simultaneously reduces the GPU memory and inference latency while achieving competitive performance. Different from cross-layer KV sharing, UniAttn does not diminish the “depth factor”. We show this in Section 3.4.
$\mathbf{x}_{i+b}^{\prime}=s_{i}V_{i+b}W_{o}+\mathbf{x}_{i+b}$ (12)
Since only the bottom block contribute to calculating the Key projection, other blocks that re-use Softmax activations do not store their corresponding K-cache, leading to GPU memory-efficiency. Moreover, removing Softmax calculations in those layers also contribute to inference latency reduction. Similar to cross-layer KV sharing, UniAttn is also orthogonal to intra-layer KV sharing. In Section 4, we employ pre-trained LLMs with GQA to evaluate UniAttn.
Linear Compensation. Simply unifying Softmax activations unavoidably introduces errors in the forward pass. Formally, the error $\epsilon$ in layer $i+1$ is:
$\epsilon$ $=\mathbf{x}_{i+b}^{\text{ori}\ \prime}-\mathbf{x}_{i+b}^{\text{uni}\ \prime}$ (13)
$=(\text{softmax}(\frac{Q_{i+b}K^{T}_{i+b}}{\sqrt{d_{k}}})-s_{i})V_{i+b}W_{o}+(\mathbf{x}_{i+b}^{\text{ori}}-\mathbf{x}_{i+b}^{\text{uni}})$
where $\mathbf{x}_{i+b}^{\text{ori}}$ and $\mathbf{x}_{i+b}^{\text{Uni}}$ denote the input to layer $i+b$ in the original model and the model with UniAttn applied, respectively. According to (Koohpayegani & Pirsiavash, 2024), calculating linear projections yield significantly lower latency compared to the Softmax operation. And thus we study the effectiveness of compensating $\epsilon$ with a linear projection matrix $W_{c}$. To compensate for $\epsilon$, we apply $W_{c}$ to each layer with unified Softmax activations as:
$\mathbf{x}_{i+b}^{\prime}=s_{i}V_{i+1}W_{o}+\mathbf{x}_{i+b}+\mathbf{x}_{i+b}W_{c}$ (14)
We initialize $W_{c}$ according to the following Theorem:
Theorem 3.7.
The optimal initialization for $W_{c}$ satisfies:
$W_{c}=V\Sigma^{+}U^{T}\mathbb{E}(\epsilon),$ (15)
We leave the proof in the Appendix. In practice, we forward a small portion of training data from the post-training datasets, and use the average values $\bar{\mathbf{x}}_{i+b}$ and $\bar{\epsilon}$ as estimations of $\mathbb{E}(\mathbf{x}_{i+b})$ and $\mathbb{E}(\epsilon)$. Considering that compensating for errors in previous layers would influence the calculation in subsequent layers, we calculate the initialization of $W_{c}$ bottom-up. Experiments on the F15K dataset (Su, 2023) with sequence length 5,120 yield errors of 11.09 (LLaMA-2 7B) and 5.76 (LLaMA-3.1 8B) after linear compensation, representing only 2.9% and 6.4% of the error before compensation, respectively, well demonstrating its effectiveness. See the appendix for more theoretical insights.
To better train the linear transformations, we adopt a two-stage post-training pipeline. In the first stage, we only fine-tune the $W_{c}$ weights and keep other weights frozen. Early stop is also adopted in the first stage. We then conduct a full fine-tuning in the second stage. We formalize a detailed pipeline in Algorithm 1 in the Appendix.
3.4 Discussion
@@
Our UniAttn reduces both memory cost and inference latency simultaneously. We underscore that unlike cross-layer KV sharing, our UniAttn does not diminish the “depth factor” $\delta$ in the forward pass.
@@
In pre-trained Pre-Norm LLMs, the “depth factor” in UniAttn forward pass is unignorable.
Proof.
In the forward pass of UniAttn:
$\mathbf{x}_{i+1}^{\prime}$ $=s_{i}(\mathbf{x}_{i+1}W_{v,i+1})W_{o}+\mathbf{x}_{i+1}$ (16)
$=s_{i}((\mathbf{x}_{i}+\delta)W_{v,i+1})W_{o}+\mathbf{x}_{i+1}$
$=s_{i}\mathbf{x}_{i}W_{v,i+1}W_{o}+s_{i}\delta W_{v,i+1}W_{o}+\mathbf{x}_{i+1}$
As $\mathbf{x}_{i}$ and $s_{i}$ have undergone the same multiplications, we can assume that UniAttn does not modify the relative size of both terms. Since we do not ignore the depth factor $\delta$, we consider $s_{i}\delta W_{v,i+1}W_{o}$ is unignorable. ∎
Table 1: Post-training performance (%) comparison. Bold and underline denote the best and second-best performance of compressed models. For each method, we report their time to first token (TTFT, in seconds) and KV-cache retain rate (KV Cache).
@@
Pre-Trained 2.26 100% 75.0 56.8 59.1 63.6 46.8 71.7 59.3
Post-Train 2.26 100% 75.4 59.0 63.2 65.9 50.1 73.9 62.0
@@ 4.1 Experimental Settings
Implementation details. We adopt the pattern in Figure 2 and apply UniAttn in the top half of layers. Unless otherwise noted, we adopt Superblock size $=4$, which yields a total of 4 Superblocks in LLaMA-2 7B, LLaMA-3.1 8B, Mistral 7B, and 5 Superblocks in Gemma-2 9B. For all experiments, we post-train LLMs for 1 epoch. We adopt all training hyperparameter values based on existing research. All training experiments were conducted on 8 H800 GPUs. For time to first token (TTFT) time measurement, we use a single H800 GPU and context length of 8,192 and measure the time to generate the first token after receiving the sequence input. See the Appendix for more details.
@@ 4.2 Performance Comparison on Post-Training
We verify the effectiveness of UniAttn on 2 open-source LLMs with different architectures, namely LLaMA-3.1 8B (with GQA) (Dubey et al., 2024) and LLaMA-2 7B (without GQA) (Touvron et al., 2023). Please refer to the Appendix for more results on Mistral 7B (Jiang et al., 2023) and Gemma-2 9B (Rivière et al., 2024). We use LLMDrop (He et al., 2024) and CLA (Brandon et al., 2024) as competing efficient architectures, as both methods exploit the cross-layer redundancies in LLMs. For LLMDrop, we bypass the MHA modules with the highest input-output similarities in the top half of the layers. For CLA, we group the top half of the layers into CLA blocks, using the same configuration as in the grouping of Superblocks. Since both competing methods drop the entire KV-cache in the operated layers, we compare two configurations: one with the same operating layers (X-Full) and another with the same KV-cache compression ratio (half operating layers, X-Half). We also conduct post-training after applying these methods.
@@
Analysis of CLA. Since CLA only reduces the latency associated with KV projections, its impact on TTFT is minimal. As a result, CLA-Half, CLA-Full, and the original model exhibit comparable TTFT times. Furthermore, CLA achieves comparable performance to LLMDrop when applied to the same number of layers. This suggests that the cross-layer attention mechanism employed by CLA is structurally equivalent to directly bypassing the MHA module. This observation supports our theoretical analysis that cross-layer KV sharing methods like CLA can diminish the depth factor in pre-trained LLMs, in turn limiting their performance.
@@ 4.3 Experimental Analysis
We choose LLaMA-3.1 8B (Dubey et al., 2024) and the medical domain to conduct experimental analysis. Please refer to the Appendix for more analysis experiments.
Table 2: Ablation study results. Bold and underline denote the best and second-best performance of compressed models.
@@
Table 3: Results of inference cost analysis. “H2O Comp.” denotes the compression rate introduced by H2O and “KV-Cache” denotes the KV-Cache retain rate.
@@
Post-Train 2.26 - 100% 65.9
CLA-Half 2.18 - 81.3% 64.9
CLA-Full 2.23 - 62.5% 63.8
@@
Figure 4: Hyperparameter analysis results (%). Left: The relationship of average accuracy and TTFT latency to the total number of grouped Superblocks (each with size 4). Right: The relationship of average accuracy and TTFT latency to the size of each Superblock (total of 12 layers that utilize unified Softmax activation).
@@
Inference Cost Analysis. We evaluate and compare inference costs to demonstrate the efficiency of UniAttn. To further maximize its potential, we integrate UniAttn with the KV-Cache compression method H2O (Zhang et al., 2023) to further reduce memory overhead by keeping only the essential and the recent tokens in the KV-cache. We report the TTFT latency and the total KV-cache retain rate. As shown in Table 3, UniAttn achieves a significant reduction in TTFT latency compared to CLA under the same KV-cache retain rate. When combined with H2O, UniAttn+H2O reduces the KV-cache retain rate to below 50% while maintaining an average performance of 65.0%, achieving both better performance, lower inference latency, and higher KV-cache compression rate than CLA. These results clearly demonstrate the inference efficiency of UniAttn, highlighting its substantial benefits for real-world deployment scenarios.
@@
Hyperparameter Analysis. We perform a hyperparameter analysis on the number and the size of Superblocks to evaluate their impact on latency and post-training performance. First, we examine the effect of varying the number of Superblocks. Using a consistent Superblock size of 4, we group different numbers of Superblocks from the top layers to the bottom layers. For instance, grouping 2 Superblocks indicates that layer 25-28 and 29-32 form the two Superblocks. The results are shown in Figure 4 (left). From the perspective of TTFT latency, increasing the number of Superblocks significantly reduces latency. From the perspective of performance, while grouping certain top layers as Superblocks does not substantially affect model performance, continuing to increase the number of Superblocks leads to sharp performance drops. Second, we analyze the impact of Superblock size. For fair comparison, we maintain a consistent total number of layers that utilize unified Softmax activations from preceding layers as 12. Since a Superblock with size $b$ unifies Softmax in $b-1$ layers, we set $b$ such that $b-1$ is a factor of 12, resulting in Superblock sizes of 2, 3, 4, 5, 7, and 13, and number of Superblocks as 12, 6, 4, 3, 2, and 1, respectively. As shown in Figure 4 (right), the size of the Superblocks does not significantly affect post-training performance. For simplicity, we use a Superblock size of 4, although fine-grained tuning of Superblock size could further improve post-training performance.

### v2

@@ 3.3 Theoretical Analysis on Existing Methods
[…]
We define $\delta$ as the “depth factor” that model applies on activations for further discussion. $\delta$ represents the extent to which model depth influences activations. Please refer to Appendix E for theoretical insights on proposing this assumption.
Analysis on Cross-layer KV Sharing: We adopt 3.2 to analyze the cross-layer KV sharing architecture, in which KV-cache is shared across multiple consecutive layers. Suppose layer $i+1$ shares the KV-cache from layer $i$, the MHA operation in layer $i+1$ can be written as:
$\mathbf{x}^{\prime}_{i+1}=\text{softmax}(\frac{QK^{T}_{i}}{\sqrt{d_{k}}})V_{i}W_{o,i+1}+\mathbf{x}_{i+1},$ (8)
@@
Please refer to Appendix G for detailed proof. Proposition 3.3 gives a rigorous proof showing that cross-layer KV sharing diminishes the impact of model depth on model activations. To offer an intuitive understanding of how KV sharing alters the structure of QKV computation within self-attention, we reformulate the operation as follows:
$\text{QKV-Comp}_{i+1}$ $=\text{softmax}(A_{i})V_{i}$ (9)
$=\text{softmax}(\frac{\mathbf{x}_{i}W_{q,i+1}K^{T}_{i}}{\sqrt{d_{k}}})V_{i}.$
Compared to QKV-compi in the previous layer, the only difference lies in the use of distinct $W_{q}$ matrices, suggesting that the QKV computation in layer $i+1$ can be viewed as introducing an additional query head applied to the same KV pairs as in layer $i$. In this sense, cross-layer KV sharing implicitly transforms QKV computation in different layers into additional query heads.
3.4 Discussion on UniAttn
@@
Our UniAttn reduces both memory cost and inference latency simultaneously. We underscore that unlike cross-layer KV sharing, our UniAttn does not diminish the “depth factor” $\delta$ in the forward pass. Please refer to Appendix H for detailed proof.
@@
In pre-trained Pre-Norm LLMs, the “depth factor” in UniAttn is unignorable.
@@
Pre-Trained 2.33 100% 75.0 56.8 59.1 63.6 46.8 71.7 59.3
Post-Train 2.33 100% 75.4 59.0 63.2 65.9 50.1 73.9 62.0
@@
Table 1: Post-training performance (%) comparison. Bold and underline denote the best and second-best performance of compressed models. We also report time to first token (TTFT) and KV-cache retain rate (KV Cache).
@@ 4.1 Experimental Settings
Implementation details. We adopt the pattern in Figure 2 and apply UniAttn in the top half of layers. Unless otherwise noted, we adopt Superblock size =4=4, which yields a total of 4 Superblocks in LLaMA-2 7B, LLaMA-3.1 8B, Mistral 7B, and 5 Superblocks in Gemma-2 9B. For all experiments, we post-train LLMs for 1 epoch (Lambert et al., 2024). We adopt all training hyperparameter values based on existing research. All training experiments were conducted on 8 H800 GPUs. For time to first token (TTFT) time measurement, we use a single H800 GPU and a context length of 8,192 and measure the time to generate the first token after receiving the sequence input. See the Appendix for more details.
@@ 4.2 Performance Comparison on Post-Training
We verify the effectiveness of UniAttn on 2 open-source LLMs with different architectures, namely LLaMA-3.1 8B (with GQA) (Dubey et al., 2024) and LLaMA-2 7B (without GQA) (Touvron et al., 2023). We use LLMDrop (He et al., 2024) and CLA (Brandon et al., 2024) as competing efficient architectures, as both methods exploit the cross-layer redundancies in LLMs. For LLMDrop, we bypass the MHA modules with the highest input-output similarities in the top half of the layers. For CLA, we group the top half of the layers into CLA blocks, using the same configuration as in the grouping of Superblocks. Since both competing methods drop the entire KV-cache in the operated layers, we compare two configurations: one with the same operating layers (X-Full) and another with the same KV-cache compression ratio (half operating layers, X-Half). We also conduct post-training after applying these methods. Same conclusions are observed on Mistral 7B (Jiang et al., 2023) and Gemma-2 9B (Rivière et al., 2024). Detailed results are in the Appendix due to the page limit.
@@
Analysis of CLA. Since CLA only reduces the latency associated with KV projections, its impact on TTFT is minimal. As a result, CLA-Half, CLA-Full, and the original model exhibit comparable TTFT times. Furthermore, CLA achieves comparable performance to LLMDrop when applied to the same number of layers. This suggests that the cross-layer attention mechanism employed by CLA is structurally equivalent to directly bypassing the MHA module. This observation supports our theoretical analysis that cross-layer KV sharing methods like CLA can diminish the depth factor in pre-trained LLMs, limiting their performance.
@@ 4.3 Experimental Analysis
We choose LLaMA-3.1 8B (Dubey et al., 2024) and the medical domain to conduct experimental analysis. See Appendix C for more analysis.
@@
Table 2: Ablation study results. Bold and underline denote the best and second-best UniAttn performance.
@@
Post-Train 2.33 - 100% 65.9
CLA-Half 2.29 - 81.3% 64.9
CLA-Full 2.22 - 62.5% 63.8
@@
Table 3: Results of inference costs analysis. “H2O Comp.” denotes the compression rate introduced by H2O and “KV-Cache” denotes the KV-Cache retain rate.
Figure 4: Comparison between baselines, UniAttn, and UniAttn without (“w/o”) linear compensation.
Figure 5: Hyperparameter analysis results on average accuracy (%) and TTFT latency (s). Left: Different total numbers of grouped Superblocks (each with size 4). Right: Different Superblock sizes (total of 12 layers that utilize unified Softmax activation).
@@
Inference Costs Analysis. We evaluate and compare inference costs to demonstrate the efficiency of UniAttn. To further maximize its potential, we integrate UniAttn with the KV-cache compression method H2O (Zhang et al., 2023) to further reduce memory overhead by keeping only the essential and the recent tokens in the KV-cache. We report the TTFT latency and the total KV-cache retain rate. As shown in Table 3, UniAttn achieves a significant reduction in TTFT latency compared to CLA under the same KV-cache retain rate. When combined with H2O, UniAttn+H2O reduces the KV-cache retain rate to below 50% while maintaining an average performance of 65.0%, achieving both better performance, lower inference latency, and higher KV-cache compression rate than CLA. These results clearly demonstrate the inference efficiency of UniAttn, highlighting its substantial benefits for real-world deployment scenarios.
Plain Softmax Unification v.s. Baselines. To further substantiate the theoretical analysis and demonstrate the effectiveness of Softmax unification, we compare UniAttn (without linear compensation, UniAttn-w/o-comp) against CLA and LLMDrop. As shown in Figure 4, simply unifying Softmax activations yields notable improvements over baselines, highlighting its core contribution. The linear projection serves as an enhancement, bringing UniAttn closer to post-train-level results.
@@
Hyperparameter Analysis. We perform a hyperparameter analysis on the number and the size of Superblocks to evaluate their impact on latency and post-training performance. First, we examine the effect of varying the number of Superblocks. Using a consistent Superblock size of 4, we group different numbers of Superblocks from the top layers to the bottom layers. For instance, grouping 2 Superblocks indicates that layer 25-28 and 29-32 form the two Superblocks. The results are shown in Figure 5 (left). From the perspective of TTFT latency, increasing the number of Superblocks significantly reduces latency. From the perspective of performance, while grouping certain top layers as Superblocks does not substantially affect model performance, continuing to increase the number of Superblocks leads to sharp performance drops. Second, we analyze the impact of Superblock size. For fair comparison, we maintain a consistent total number of layers that utilize unified Softmax activations from preceding layers as 12. Since a Superblock with size $b$ unifies Softmax in b1b-1 layers, we set $b$ such that b1b-1 is a factor of 12, resulting in Superblock sizes of 2, 3, 4, 5, 7, and 13, and number of Superblocks as 12, 6, 4, 3, 2, and 1, respectively. As shown in Figure 5 (right), the size of the Superblocks does not significantly affect post-training performance. For simplicity, we use a Superblock size of 4, although fine-grained tuning of Superblock size could further improve post-training performance.
Figure 6: Generation efficiency (wall-clock time seconds) comparison on LLaMA-3.1 8B with 8192 context length.
Method Sequence Length AVG
4096 8192 16384 32768
Pre-Trained 82.67 0 0 0 20.67
Post-Train 86.67 87 80.33 72 81.5
LLMDrop-Half 87 83.67 77.67 68.67 79.25
CLA-Half 86.67 85.67 78.67 69.33 80.08
UniAttn (Ours) 87 86.67 79.67 70.33 80.92
Table 4: Post-training performance (%) comparison on Niah Multikey in RULER (Hsieh et al., 2024). Bold and underline denote the best and second-best performance.
Method Grouping PubMed MedMCQA MedQA AVG
Post-Train - 75.4 59.0 63.2 65.9
UniAttn Fixed size (4) 79.0 57.6 59.5 65.4
Adaptive 78.8 57.3 59.5 65.2
Table 5: Performance (%) comparison of different Superblock grouping methods in UniAttn. Bold denotes the best performance.
Generation Efficiency Analysis. To further analyze inference efficiency, we present the generation results of UniAttn against the comparing baselines, namely CLA-Half, CLA-Full, LLMDrop-Half, and LLMDrop-Full. We measure the end-to-end generation time with the LLaMA-3.1 8B backbone, and with generating 1024, 2048, and 4096 tokens using a context length of 8192. As shown in Figure 6, across all comparing methods, UniAttn consistently achieves Top-2 latency, only slightly behind LLMDrop-Full where all operated layers are simply skipped. These results demonstrate that unifying Softmax operations can substantially reduce the inference overhead during generation.
Generalization to Long-Context Modeling. To further validate the effectiveness of UniAttn on long context modeling scenarios, we follow the training recipe of (Bai et al., 2024) and post-train LLaMA-2 7B, which lacks native support for extended contexts. We compare UniAttn with two strong baselines, CLA-Half and LLMDrop-Half, on the Niah Multikey task in the RULER benchmark (Hsieh et al., 2024). To successfully implement long-context modeling, here we conduct experiments with FlashAttention (Dao et al., 2022) applied. As shown in Table 4, UniAttn consistently ranks in the top two (second only to full post-training) and outperforms all competing methods, well demonstrating its strong generalization ability to complex long-context tasks. This also suggests that our UniAttn is compatible with common acceleration frameworks such as FlashAttention.
Adaptive Superblock Grouping Schedule. To further investigate layerwise redundancy–aware grouping, we performed an additional experiment using a similarity-based adaptive scheme. We controlled the total number of “grouped layers” and assigned Superblocks by cosine similarity of layerwise Softmax activations. As shown in Table 5, for LLaMA-3.1 8B on the medical domain, this similarity-based grouping variant achieved comparable performance to setting the Superblock size as a fixed number, indicating that the performance of UniAttn is robust to different grouping strategies, and that adaptive grouping does not yield meaningful gains over the simpler fixed-size design.